OpenStudy (anonymous):

what is the value of f(x)=2(0.75)^x when x=3?

OpenStudy (mathmale):

First, some preliminaries: What is 2^3? What is 10^3? what is 0.75^3?

OpenStudy (anonymous):

2^3=8 10^3= 1000 0.75^3= .42

OpenStudy (mathmale):

Great. Look at the original problem. Note how I let x=3. What is 2(2)^3? what is 2(10)^3? What is 2(0.75)^3? Make use of your previous results. No need to start over again.

OpenStudy (anonymous):

2(2)^3= 16 2(10)^3 = `2000 2(0.75) ^3 = .84

OpenStudy (anonymous):

where did we get 10 from?